OpenStudy (anonymous):

PLEASE HELP ME d + e = 1 −d + e = −5 Label the ordered pair as (d, e). (0, 0) (3, −2) (−2, 3) (−3, 0)

OpenStudy (michele_laino):

if we add tose equations, together, we get: \[\left( {d + e} \right) + \left( { - d + e} \right) = 1 + \left( { - 5} \right)\] please simplify

OpenStudy (michele_laino):

those*

OpenStudy (anonymous):

ok so is it -3,2

OpenStudy (michele_laino):

no, since if we simplify that expression above, we get: \[\begin{gathered} 2e = - 4 \hfill \\ e = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

-2

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

now, please substitute e=-2, into the first equation, namley: \[d + e = 1\]

OpenStudy (michele_laino):

hint: we get: \[d + \left( { - 2} \right) = 1\] what is d?

OpenStudy (anonymous):

b